OpenStudy (anonymous):

Calculate the measure of each angle of an octagon, where the measures of an angle form an arithmetic sequence and the least measure is 58 degrees.

OpenStudy (campbell_st):

the angle sum of a ploygon is (n - 2)*180 = 1080 if its a arithmetic sequence you know the 1st term a = 58, n = 8 you need to find the common difference Use the sum of an arithmetic sequence to find the common difference \[1080 = \frac{8}{2}[2\times58 + ( 8 - 1)d]\] solve for d \[\frac{2160}{8} = 116 + 7d\] \[270 = 116 + 7d\] d = 22 the common difference is 22 degrees then the angles are 58, 80, 102, 124, 146, 168, 190, 212